The following function is cumulative distribution function. 5 7 %520 T<-15 0.15-15 r<25 (r)一亻 0.70 25 、〈45 < 45X -

(a) Find P(X ≤ 45).


(b) Find P(35 ≤ X ≤ 55).

(c) Find the probability mass function of X.
Answer #1

a)P(X<=45)=F(45)=1

b)P(35<=X<=55)=F(55)-F(35)=1-0.7 =0.3

c) probability mass function of X is as follows:

P(X=-15)=0.15

P(X=25)=0.7-0.15=0.55

P(X=45)=1-0.7=0.3